OpenStudy (anonymous):

What is the equation in standard form of a parabola that contains the following points? (-2, -20) (0, -4) (4, -20) Please help, explanation would be much appreciated

OpenStudy (nikato):

ok, do you know what kind of equation makes a parabola?

OpenStudy (anonymous):

y = a(h-x)^2 + k where x and y are points on the parablola [given] h and k is the vertext point of a parabola [given] a is a 'multiplier' or a 'scalar' for how wide the parabola spans out [which you'll need to solve for] make sense so far?

OpenStudy (nikato):

or you can just substittue x and y in a quadratic equation, y=Ax^2 + Bx + C, and use system of equaions

OpenStudy (anonymous):

Okay, I think I get it so far

OpenStudy (nikato):

whose method do you wanna use?

OpenStudy (anonymous):

We can use yours

OpenStudy (anonymous):

:( haha ^_^

OpenStudy (nikato):

HAHA^ so can you substitute x and y in that equation i gave you. i'll start with the first one. you try the other two -20= A(-2)^2 + B(-2) + C

OpenStudy (anonymous):

-4 = a(0)^2 + b(0) + c ??

OpenStudy (nikato):

yes. and the last one?

OpenStudy (anonymous):

-20 = a(4) + b(4) + c

OpenStudy (nikato):

 ok, do you see that these are the 3 equations we came up with?